TASK 2 - PC 4.2 Calculate material mix and yield variances
Using the following equations for Material or Quantity Variance Model.
Note: SP – Standard Price; SQ – Standard Quantity; AQ – Actual Quantity; AP – Actual Price
SP (AQ – SQ) and AQ (AP – SP) calculate the following variances using the question below:

Kaltham LLC manufactures zips for the fabric industry in the UAE, she has the following direct material standard to manufacture one zip.
1.5 pounds per zip at Dh 4.20 per pound
Last 2 weeks 3,400 pounds of materials were bought from the supplier and used to make 2,000 zips for a customer in Abu Dhabi. The total cost of the materials was Dh 13,770.


a)   Calculate the actual price per zip:


b)   Calculate the price variance:

c)   Calculate the quantity variance and indicate whether is adverse or favourable:
Answer #1

a).  Actual price per zip

Total cost of materials = Dh 13,770

Actual output = 2,000 zips

Actual price per zip = Total cost of materials / Actual output

=13,770/2,000

=6.885 per zip

So, Actual price per zip is 6.885

b). Price variance

Price variance is computed by subtracting standard price from actual price then multiplying with atual quantity

Price variance = (actual price - standard price ) X Actual quantity

First , we need to compute actual price of the material

We have used 3,400 pounds of material to produce actual output i.e 2,000 zips

Total cost of material = Dh 13,770

Actual price of the material = Total cost of the material / Actual quantity used to produce the output

=13,770/3,400

=Dh 4.05

Now, we have all the Required information to compute the price variance

Actual price = Dh 4.05 per pound

Standard price = Dh 4.20 per pound

Actual quantity used to produce the output = 3,400 pounds

Price variance = (actual price - standard price ) X Actual quantity

=(4.05 - 4.20) X 3,400

= -Dh 510 ( favourable)

So, the pice variance is -Dh 510

As in this case, actual price is less than standard price so, it is favourable

c). Quantity variance

Quantity variance is computed by subtracting standard quantity for actual output from actual quantity then, multiplying with standard price

Quantity variance = (Actual quantity used - Standard quantity for actual output ) X standard price

First of all , we need to compute standard quantity for actual output

Here, standard requirement is given which is 1.5 pounds per zip and actual output is given as 2,000 zips

Standard quantity for actual output = standard requirement X actual output

= 1.5 X 2,000

=3,000 pounds

Now, we have all required information

Standard quantity for actual output = 3,000 pounds

Actual quantity used = 3,400 pounds

Standard price = Dh 4.20 per pound

Quantity variance = (Actual quantity used - Standard quantity for actual output ) X standard price

= (3,400-3,000) X 4.20

= Dh 1,680

So, Quantity variance is Dh 1,680

As , the actual quantity used is more than the standard quantity for actual output , so it is adverse